OpenStudy (anonymous):

Write 5.925 as a mixed number in simplest form.

OpenStudy (mridul):

237/40

OpenStudy (anonymous):

Mridul, is supposed to be simplest form

OpenStudy (diyadiya):

5925/1000 =237/40

OpenStudy (mridul):

how does it get more simpler???

OpenStudy (diyadiya):

Now you cant reduce further

OpenStudy (anonymous):

I do my math online and it gets atomaticllly checked...I got it wrong

OpenStudy (anonymous):

it says it was: 5 37/40

OpenStudy (anonymous):

I am confused

OpenStudy (diyadiya):

ok Yeah that is in mixed fraction

OpenStudy (mridul):

check d q. again

OpenStudy (anonymous):

This is there explenation: Write the decimal as a mixed number with 1,000 as the denominator. Reduce the mixed number to simplest form. 5.925 = 5 925/ 1,000 = 5 + 925/ 1,000 = 5 + 925 ÷ 25 1,000 ÷ 25 = 5 + 37/40 = 5 37/40

OpenStudy (anonymous):

Don't know

OpenStudy (diyadiya):

\[5 \frac{37}{40} = \frac{5 \times 40 + 37}{40} =\frac{200 +37}{40} = \frac{237}{40}\]

OpenStudy (anonymous):

hmmm ok well thanks guys anyways

OpenStudy (anonymous):

I apprepcatie your help

OpenStudy (mridul):

welcome

OpenStudy (diyadiya):

Divide 237 /40 using long Division and you can write it in mixed fraction as quotient + remainder/divisor

OpenStudy (anonymous):

ok thanks alot Diyadiya